OpenStudy (anonymous):

Alg 2 help, Fan + Medal

OpenStudy (anonymous):

A group of adults and students went on a class trip to Washington, DC. The number of male students was 1 more than 7 times the number of adults. The number of female students was half the number of male students. If the total number of people who went on the trip is 82, find the numbers of male students and female students.

OpenStudy (anonymous):

The number of male students is...

OpenStudy (anonymous):

A. 36 B. 50 C. 70

OpenStudy (anonymous):

The number of female students is...

OpenStudy (anonymous):

D. 18 E. 25 F. 35

OpenStudy (the_fizicx99):

Your attempt?

OpenStudy (anonymous):

I got B and E

OpenStudy (anonymous):

That's the closest

OpenStudy (the_fizicx99):

My question is \(\ how\) you got those answers.

OpenStudy (anonymous):

x+7x+1+(7x+1/2)=82

OpenStudy (anonymous):

@mustafa2014 am I right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

x=number of male male+female+adults=82 x+(x-1)/6+0.5x=82 x=50

OpenStudy (anonymous):

srry x+(x-1)/7+0.5x=82

OpenStudy (anonymous):

7 not 6